OpenStudy (anonymous):

Determine if (3/4 x -2) is a solution to -4x - 7y=11.

OpenStudy (earthcitizen):

is this all ?

OpenStudy (anonymous):

yes this is the whole problem

OpenStudy (earthcitizen):

hmmph..i'm assuming y=3/4x-2, then just plug that into the second equation

OpenStudy (anonymous):

i just really dont get it at all

OpenStudy (mertsj):

I don't think you posted the problem right.

OpenStudy (mertsj):

Ordered pairs are solutions to equations with two variables. Is there a comma between the 3/4 and x-2?

OpenStudy (anonymous):

im sorry here it is correctly... (3/4 times negative 2) is it a solution to _4x-7y=11

OpenStudy (anonymous):

i mean -4

OpenStudy (anonymous):

on second part

OpenStudy (mertsj):

It doesn't make sense. It should say something like: Is (5,3) a solution to -4x -7y = 11

OpenStudy (mertsj):

Would you please try really hard to type it exactly like it is in the book?

OpenStudy (anonymous):

ok like seriously im doing graphs like x an y coordinate type thing...idunderstand either but here goes it says determine if (3 over 4 then theres a dot -2) is a solution to -4x-7y=11. show work this is my nquestion as it appears

OpenStudy (mertsj):

Good. Now that makes sense.

OpenStudy (mertsj):

It's not a dot, it's a comma.

OpenStudy (anonymous):

ok but it doesnt look like one it in the middle between problems so i thought it meant times or multiply sorry :)

OpenStudy (mertsj):